OpenStudy (jabez1777):

My teacher went over this with us this morning, may someone help me again? :)

OpenStudy (jabez1777):

http://prntscr.com/dcym4e

OpenStudy (mirving):

i will <3 2 help u

OpenStudy (daniyil_the_spy):

ok go for it

OpenStudy (mirving):

d?

OpenStudy (dastudyguy12):

D

OpenStudy (jabez1777):

No direct answers, please. :(

OpenStudy (jabez1777):

@triciaal, may you help me? :)

OpenStudy (sshayer):

\[ 4^{-(y+1)}=64=4^3\] -(y+1)=3 ?

OpenStudy (triciaal):

are you learning logs? or just exponents 64 can be expressed as 4^3 rules for exponents

OpenStudy (sshayer):

\[\left( \frac{ 1 }{ 4 } \right)^{y+1}=\left(1* 4^{-1} \right)^{y+1}=\left\{ \left( 4 \right)^{-1} \right\}^{y+1}=4^{-(y+1)}\]

OpenStudy (jabez1777):

Just exponents hahaha!! Tkhunny was using Logarithms earlier that's what confused me. :P

OpenStudy (triciaal):